Calculate the resistance of the resistor required by a galvanometer of resistance 50 ohms which gives a full scale deflection for 1mA, to be adapted to measure currents of up to 5A
1
Expert's answer
2019-08-26T10:56:03-0400

The resistance of the shunt


"R_{SH}=\frac{I_G}{I-I_G}R_G=\frac{R_G}{I/I_G-1}"

"R_{SH}=\frac{50\:\Omega}{5/0.001-1}=0.01\:\Omega"